# Multi-Modal Federated Learning for Cancer Staging over Non-IID Datasets with Unbalanced Modalities

### v1

@@ Abstract
The use of machine learning (ML) for cancer staging through medical image analysis has gained substantial interest across medical disciplines. When accompanied by the innovative federated learning (FL) framework, ML techniques can further overcome privacy concerns related to patient data exposure. Given the frequent presence of diverse data modalities within patient records, leveraging FL in a multi-modal learning framework holds considerable promise for cancer staging. However, existing works on multi-modal FL often presume that all data-collecting institutions have access to all data modalities. This oversimplified approach neglects institutions that have access to only a portion of data modalities within the system. In this work, we introduce a novel FL architecture designed to accommodate not only the heterogeneity of data samples, but also the inherent heterogeneity/non-uniformity of data modalities across institutions. We shed light on the challenges associated with varying convergence speeds observed across different data modalities within our FL system. Subsequently, we propose a solution to tackle these challenges by devising a distributed gradient blending and proximity-aware client weighting strategy tailored for multi-modal FL. To show the superiority of our method, we conduct experiments using The Cancer Genome Atlas program (TCGA) datalake considering different cancer types and three modalities of data: mRNA sequences, histopathological image data, and clinical information.
@@ I Introduction
In this work, we take an initial step towards implementing multi-modal FL in a scenario where clients not only have non-iid (non-independent and identically distributed) data, but also possess an uneven number of data modalities. We present a system model that eliminates the need for clients to possess identical sets of data modalities by employing a versatile distributed encoder-decoder architecture. Within this framework, the encoders are tasked with the extraction of features from each distinct data modality at each client. These encoders and decoders/classifiers will undergo separate aggregations in order to address the heterogeneity of modalities across clients. This approach ensures that all trainable components relevant to a specific data modality are collectively trained using the data of that particular modality across all clients. The final global model resulting from this training can then be applied over the combination of all data modalities.
We show that while our proposed system model holds considerable promise, it confronts a series of challenges. A notable concern is the disparity in the training and convergence speeds of various data modalities among different clients. This variability can lead to certain encoders rapidly exhibiting a biased behavior towards local data of clients, which can impede the global model’s convergence. [^1] This phenomenon, in turn, may result in overfitting toward specific data labels, particularly in the context of generic FL classification tasks, due to the imbalanced effect of each modality during local ML training. To tackle this challenge, we introduce a novel approach termed distributed gradient blending (DGB) for FL. Although conceptually distinct, this approach draws inspiration from a seminal work on centralized multi-model ML [21]. Our proposed DGB method capitalizes on information pertaining to the performance of participating clients in FL and the characteristics of their respective modalities to determine effective back-propagation weights for local training of each encoder at each client. To further enhance the robustness of our proposed method against non-iid data distribution across the clients, we propose proximity-aware client weighting (PCW). PCW considers the quality of data at clients to further enhance the parameters utilized in DGB.
@@
  • We conduct numerical evaluations on a dataset of three cohorts gathered by The Cancer Genome Atlas program (TCGA - https://www.cancer.gov/tcga) for three different cancer types (breast, lung, and liver) and demonstrate that our method can lead to significant performance gains as compared to various baseline methods in multi-modal FL.
@@ III Network and Learning Model
In this section, we first present our network model of interest (Sec. III-A). Subsequently, we offer a concise discussion on conventional uni-modal FL (Sec. III-B). Our contribution unfolds as we extend the conventional FL paradigm by introducing a novel FL system designed to accommodate a network of clients that possess an uneven distribution/number of data modalities (Sec. III-C). We conclude this section by providing an overview of the pivotal challenges that necessitate addressing to achieve high-performance ML models within the context of our proposed FL system (Sec. III-D).
@@ III-B Conventional Uni-modal Federated Learning
In the conventional uni-modal FL framework, the datasets held by institutions, and consequently the global dataset, consist solely of a single data modality (e.g., image data) [13, 26, 27]. FL conducts ML model training through a series of local (model) training rounds indexed by $k\in\in\{0,\cdots,K-1\}$, each leading to a global (model) aggregation step, indexed by $t\in\{0,\cdots,T-1\}$. Within this context, each institution $n$ trains a local ML model, represented by a parameter vector $\bm{\omega}_{n}$. The initial local training round commences with the server initializing a global model $\bm{\omega}^{(t)}$ at $t=0$ and disseminating it among the institutions. Subsequently, each institution $n$ synchronizes its local model as $\bm{\omega}_{n}\leftarrow\bm{\omega}^{(t)}$ for $t=0$ and utilizes it as the initial model for local training, i.e., $\bm{\omega}^{(t),0}_{n}\leftarrow\bm{\omega}_{n}$. Following this initialization, each client undergoes $K$ local stochastic gradient descent (SGD) iterations, utilizing its own dataset to obtain $\bm{\omega}^{(t),K}_{n}$. The interim ML updates, for $10\leq k\leq K-$ 1, are computed as $\bm{\omega}^{(t),k+1}_{n}=\bm{\omega}^{(t),k}_{n}-\eta^{(t),k}\widetilde{\nabla}\mathcal{L}_{n}(\bm{\omega}^{(t),k}_{n})$, where $\eta^{(t),k}$ denotes the step-size/learning-rate, and $\mathcal{L}_{n}(\cdot)$ denotes the local loss function of the institution $n$, which quantifies the predictive accuracy of the ML model over its local dataset. Further, ~$\widetilde{\nabla}\mathcal{L}_{n}(\cdot)$ denotes the stochastic gradient obtained by sampling a random mini-batch of data from the local dataset $\mathcal{D}^{\mathsf{Tr}}_{n}$.
@@
This global model is broadcast to institution, using which they synchronize their local models as mentioned above and initiate the next round of local ML training and global aggregation.
@@ III-C1 Multi-modal FL under Unbalanced Modalities
In real-world scenarios, institutions may have access to various subsets of modalities. This phenomenon makes the conventional multi-modal FL, which is based on naïve extension of uni-modal FL to multi-modal settings [31, 32, 20] impractical, as these works presume the existence of the same set of data modalities across all the institutions. In particular, in a network of institutions with non-uniform number of data modalities, the number of encoders used across them will differ. Further, the size of the input feature fed to the classifier will also vary across institutions. These variabilities make the naïve ML training and aggregation methods of [31, 32, 20] ineffective. Consequently, we propose a generalized distributed encoder-decoder architecture, which increases the flexibility of model aggregations at the server by aggregating (i) each modality’s encoder and (ii) classifiers of different existing modality combinations across the institutions separately. A schematic of our proposed system and model aggregation strategy is depicted in Fig. 1(a) and 1(b), respectively.
@@
Let $\mathcal{M}$ and $\mathcal{Y}$ denote the sets encompassing all data modalities within our network distributed across the institutions, and the data labels (i.e., cancer stages in our scenario), respectively. Moreover, let $\mathcal{M}_{n}\subseteq\mathcal{M}$ denote the subset of modalities held by institution $n$, which are present in its local dataset $\mathcal{D}_{n}$. Considering each institution $n$, each data point $d\triangleq(\bm{X},\bm{y})\in\mathcal{D}_{n}$ is associated with a feature vector $\bm{X}$ and a label $\bm{y}$, where $\bm{X}=\{\bm{x}_{m}\}_{m\in\mathcal{M}_{n}}$ contains the set of features corresponding to each modality $m$ possessed by the institution. Additionally, $\bm{y}\in\mathcal{Y}$ is represented in a one-hot encoded format, capturing the cancer stage. As will be formalized next, to handle clients with an uneven distribution of modalities, we will put a greater emphasis on improving the local training of encoders. This approach diverges from the recent methodologies in multi-modal FL [32], where the emphasis lies on aggregating the classification layers – and co-attention layers – which adapt well to scenarios where clients have access to all the modalities. This paradigm shift allows our approach to extend the distributed ML training framework of FL for each modality to all institutions possessing data pertaining to that specific modality. In particular, in our approach, the encoder used for each modality would benefit from the commonality of data across the institutions possessing that modality. This, in turn, eliminates the restricting assumption that institutions must possess all data modalities.
@@
where $\bm{\omega}_{m_{i},n}(\mathbf{x}_{m_{i}})$, for each $i\in\{1,\cdots,|\mathcal{M}_{n}|\}$, represent the encoder output upon being fed with data feature of modality $m_{i}$. Also, the function $\bm{\varpi}_{n}(\cdot)$ encompasses the description of the decoder/classifier, which takes as its input the concatenated/fused feature of all the modalities available at the institution and produces a predicted label as its output.
Fig. 2: A schematic of the encoder-decoder structure at an arbitrary institution $n$ with three data modalities.
@@ III-C2 Machine Learning Task Formulation
We consider the loss function at institution $n$ for an arbitrary local model parameter realization $\bm{\omega}_{n}$ as $\mathcal{L}_{n}(\bm{\omega}_{n};\mathcal{D}_{n})=\frac{1}{|\mathcal{D}_{n}|}\sum_{d\in\mathcal{D}_{n}}\ell(\bm{\omega}_{n};d)$ where $\ell(\bm{\omega}_{n};d)$ is the ML model loss function evaluated on datapoint $d$. In this work, we use cross-entropy to measure the loss [34]. Also, we use $\bm{\omega}_{n}\triangleq\left[\{\bm{\omega}_{m,n}\}_{m\in\mathcal{M}_{n}}:\bm{\varpi}_{n}\right]$ to denote the concatenation of the encoders model parameters in conjunction with the decoder for institution $n$.
@@
$\mathcal{P}(\mathcal{M})\triangleq\{\mathcal{M}_{n}\}_{n\in\mathcal{N}}.$ (2)
If all the institutions have the same modality ($\mathcal{M}_{n}=\mathcal{M},\forall n$), then $\mathcal{P}(\mathcal{M})=\mathcal{M}$. Subsequently, since institutions with the same combinations of modalities use the same architecture of classifiers, we denote $\widehat{\bm{\varpi}}_{C}$ as the common architecture of the classifier used in the system, for all institutions $n$ with modality combination $C\in\mathcal{P}(\mathcal{M})$, for which $|\widehat{\bm{\varpi}}_{C}|=|\bm{\varpi}_{n}|$, $\forall n$, where |.||.|| . | captures the number of classifier parameters. [^4] For example, if the system has one institution with modality $m_{1}$, one institution with modality $m_{2}$, one institution with modality $m_{3}$, one institution with modality $m_{1},m_{2}$, and one institution with modality $m_{1},m_{2},m_{3}$, then, $\mathcal{P}(\mathcal{M})=\{\{m_{1}\},\{m_{2}\},\{m_{3}\},\{m_{1},m_{2}\},\{m_{1},m_{2},m_{3}\}\}$. In this case, $C$ can be any of the five sets that constitute $\mathcal{P}(\mathcal{M})$.
@@
$\mathcal{L}(\bm{\omega})\triangleq\sum_{n\in\mathcal{N}}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\mathcal{L}_{n}(\bm{\omega}_{n}).$ (3)
Therefore, our goal is to find a set of global model parameters, consisting of a set of model parameters for encoders used for different modalities and the decoders. Mathematically, we are interested in obtaining $\bm{\omega}^{\star}$, which is given by
$\bm{\omega}^{\star}=\big{\{}\{\widehat{\bm{\omega}}^{\star}_{m}\}_{m\in\mathcal{M}},\{\widehat{\bm{\varpi}}_{C}^{\star}\}_{C\in\mathcal{P}(\mathcal{M})}\big{\}}\triangleq\operatorname*{arg\,min}_{\bm{\omega}\in\mathbb{R}^{|\bm{\omega}|}}\mathcal{L}(\bm{\omega}),$ (4)
where $|\bm{\omega}|=\sum_{m\in\mathcal{M}}|\widehat{\bm{\omega}}_{m}|+\sum_{C\in\mathcal{P}(\mathcal{M})}|\widehat{\bm{\varpi}}_{C}|$.

### v2

@@ Abstract
The use of machine learning (ML) for cancer staging through medical image analysis has gained substantial interest across medical disciplines. When accompanied by the innovative federated learning (FL) framework, ML techniques can further overcome privacy concerns related to patient data exposure. Given the frequent presence of diverse data modalities within patient records, leveraging FL in a multi-modal learning framework holds considerable promise for cancer staging. However, existing works on multi-modal FL often presume that all data-collecting institutions have access to all data modalities. This oversimplified approach neglects institutions that have access to only a portion of data modalities within the system. In this work, we introduce a novel FL architecture designed to accommodate not only the heterogeneity of data samples, but also the inherent heterogeneity/non-uniformity of data modalities across institutions. We shed light on the challenges associated with varying convergence speeds observed across different data modalities within our FL system. Subsequently, we propose a solution to tackle these challenges by devising a distributed gradient blending and proximity-aware client weighting strategy tailored for multi-modal FL. To show the superiority of our method, we conduct experiments using The Cancer Genome Atlas program (TCGA) datalake considering different cancer types and three modalities of data: mRNA sequences, histopathological image data, and clinical information. Our results further unveil the impact and severity of class-based vs type-based heterogeneity across institutions on the model performance, which widens the perspective to the notion of data heterogeneity in multi-modal FL literature.
@@ I Introduction
In this work, we take an initial step towards implementing multi-modal FL in a scenario where clients not only have non-iid (non-independent and identically distributed) data, but also possess an uneven number of data modalities. We present a system model that eliminates the need for clients to possess identical sets of data modalities by employing a versatile distributed encoder-classifier architecture. Within this framework, the encoders are tasked with the extraction of features from each distinct data modality at each client. These encoders and classifiers/decoders will undergo separate aggregations in order to address the heterogeneity of modalities across clients. This approach ensures that all trainable components relevant to a specific data modality are collectively trained using the data of that particular modality across all clients. The final global model resulting from this training can then be applied over the combination of all data modalities.
We show that while our proposed system model holds considerable promise, it confronts a series of challenges. A notable concern is the disparity in the training and convergence speeds of various data modalities among different clients. As we will later show (Sec. III-D), based on the combination of the modalities available in a specific client’s data, each modality can converge at a different speed. This variability can lead to certain encoders rapidly exhibiting a biased behavior towards local data of clients, which can impede the global model’s convergence. [^1] This phenomenon, in turn, may result in overfitting toward specific data labels, particularly in the context of generic FL classification tasks, due to the imbalanced effect of each modality during local ML training. To tackle this challenge, we introduce a novel approach termed distributed gradient blending (DGB) for FL. Although conceptually distinct, this approach draws inspiration from a seminal work on centralized multi-model ML [21]. In a nutshell, DGB capitalizes on information pertaining to the performance of participating clients in FL and the characteristics of their held data modalities to determine effective back-propagation weights for local training of each encoder-classifier pair at clients. In particular, in DGB, the back-propagation weights (i.e., the learning rate that tunes the sensitivity to the obtained gradients during the training process) is tuned based on the growth in the performance of the modalities in terms of their generalization and overfitting behavior. Nevertheless, DGB by itself cannot take into account for the heterogeneity (i.e., non-iid-ness) of data distributions [^2] In this paper, heterogeneity of data distribution is captured via both class-based heterogeneity (e.g., each client can have different ratio of data points belonging to different stages of cancer, such as stage I and II, in its dataset) and type-based heterogeneity (e.g., clients possessing data of various cancer types, such as breast, lung, and liver cancers). across clients, which is a unique challenge in distributed networks of interest to FL. Subsequently, to enhance the robustness of our proposed method against non-iid data distribution across the clients, we propose proximity-aware client weighting (PCW). In a nutshell, PCW further tunes the back-propagation weights in DGB via explicitly taking into account for the quality of data at each client, measured via the similarity of the cumulative gradient of the local ML model of the client and that of the global model.
@@
  • We conduct numerical evaluations on a dataset of three cohorts gathered by The Cancer Genome Atlas program (TCGA - https://www.cancer.gov/tcga) for three different cancer types (breast, lung, and liver) and demonstrate that our method can lead to significant performance gains as compared to various baseline methods in multi-modal FL. In our experiments, we further study the impact of class-based vs type-based heterogeneity across institutions on the model training performance. This study opens a new perspective to the notion of data heterogeneity not only in FL literature focused on downstream task of cancer staging but also in the broader multi-modal FL literature.
@@ III Network and Learning Model
In this section, we first present our network model of interest (Sec. III-A). Subsequently, we offer a concise discussion on conventional uni-modal FL (Sec. III-B). Our contribution unfolds as we extend the conventional FL paradigm by introducing a novel FL system designed to accommodate a network of clients that possess an uneven distribution/number of data modalities (Sec. III-C). We conclude this section by providing an overview of the pivotal challenges that necessitate addressing to achieve high-performance ML models within the context of our proposed FL system (Sec. III-D). The notations used in the modeling are summarized in Table I.
@@ III-B Conventional Uni-modal Federated Learning
In the conventional uni-modal FL framework, the datasets held by institutions, and consequently the global dataset, consist solely of a single data modality (e.g., image data) [13, 26, 27]. FL conducts ML model training through a series of local (model) training rounds indexed by $k\in\{0,\cdots,K-1\}$, each leading to a global (model) aggregation step, indexed by $t\in\{0,\cdots,T-1\}$. Within this context, each institution $n$ at global aggregation step $t$ and local training round $k$ trains a local ML model, represented by a parameter vector $\bm{\omega}_{n}^{(t),k}$. The initial local training round commences with the server initializing a global model $\bm{\omega}^{(t)}$ at $t=0$ and disseminating it among the institutions. Subsequently, each institution $n$ synchronizes its local model as $\bm{\omega}_{n}^{(t),k}\leftarrow\bm{\omega}^{(t)}$ for $t=0$ and $k=0$ and utilizes it as the initial model for local training, i.e., $\bm{\omega}^{(0),0}_{n}\leavevmode\nobreak\ \leftarrow\leavevmode\nobreak\ \bm{\omega}^{(0)}$. Following this initialization, each institution undergoes $K$ local stochastic gradient descent (SGD) iterations, utilizing its own dataset to obtain $\bm{\omega}^{(t),K}_{n}$. The interim ML updates, for $10\leq k\leq K-$ 1, are computed as $\bm{\omega}^{(t),k+1}_{n}=\bm{\omega}^{(t),k}_{n}-\eta^{(t),k}\widetilde{\nabla}\mathcal{L}_{n}(\bm{\omega}^{(t),k}_{n})$, where $\eta^{(t),k}$ denotes the step-size/learning-rate, and $\mathcal{L}_{n}(\cdot)$ denotes the local loss function of the institution $n$, which quantifies the predictive accuracy of the ML model over its local dataset. Further, ~$\widetilde{\nabla}\mathcal{L}_{n}(\cdot)$ denotes the stochastic gradient obtained by sampling a random mini-batch of data from the local dataset $\mathcal{D}^{\mathsf{Tr}}_{n}$.
TABLE I: Major notations used in the paper.
Notation Description
$\mathcal{N}$ Set of participating institutions
$n$ Sample institution index
$\mathcal{D}_{n}$ Set of local datapoints of institution $n$
$\mathcal{D}_{n}^{\mathsf{Tr}}$ Set of local train datapoints of institution $n$
$\mathcal{D}_{n}^{\mathsf{Va}}$ Set of local validation datapoints of institution $n$
$\mathcal{D}$ Global dataset
$d$ Sample datapoint consisting of features and class label
$\bm{X}$ Set of input features of an arbitrary datapoint
$\bm{y}$ Class label of an arbitrary datapoint
$\bm{x}_{m}$ Input features for modality $m$ of a datapoint $d=(\bm{X}$, $\bm{y})$
$K$ Number of local training rounds across institutions
$k$ Local training round index
$T$ Number of global aggregation steps
$t$ Global aggregation step index
$\bm{\omega}_{n}^{(t),k}$ Local ML model parameters for institution $n$ at global aggregation step $t$ and local training round $k$
$\bm{\omega}_{m,n}^{(t),k}$ Institution $n$’s encoder parameters for modality $m$ at global aggregation $t$ and local round $k$
$\bm{\varpi}_{n}^{(t),k}$ Classifier parameters for institution $n$’s ML model
$\widehat{\bm{\omega}}_{m}^{(t)}$ Global encoder of modality $m$ at global aggregation step $t$
$\widehat{\bm{\varpi}}_{C}^{(t)}$ Global classifier of modality combination $C$ at global aggregation step $t$
$\bm{\omega}^{(t)}$ Global model parameters at global aggregation step $t$
$\bm{\omega}_{\mathsf{obj}}$ Set of all the encoder and classifier parameters across the institutions
$\bm{\omega}_{\mathsf{obj}}^{\star}$ Optimal set of ML model weights for all institutions
$\mathcal{B}_{n}^{(t),k}$ Stochasticly selected mini-batch of data from institution $n$’s local dataset at global aggregation step $t$ and local training round $k$
$\eta^{(t),k}$ Learning rate at global aggregation $t$ and local round $k$
~$\widetilde{\nabla}\mathcal{L}_{n}$ Stochastic gradient of institution $n$
$\mathcal{M}$ Set of all available modalities across the institutions
$\mathcal{M}_{n}$ Data modality combination held by institution $n$
$P(\mathcal{M})$ Set of all existing modality combinations in the network
$C$ An arbitrary combination of modalities
$\overline{\rho}_{C,n}^{(t)}$ Normalized client weight for institution $n$ with modality combination $C$ at global aggregation step $t$
$\mathcal{L}_{n}^{\mathsf{Va}}$ Validation loss of institution $n$
$\mathcal{L}_{n}^{\mathsf{Tr}}$ Training loss of institution $n$
$O_{C}^{(t)}$ Overfitting for modality combination $C$ at global aggregation step $t$
$G_{C}^{(t)}$ Generalization for modality combination $C$ at global aggregation step $t$
$\Gamma_{m,n}^{(t)}$ DOGR coefficent at global aggregation step $t$ for the encoder of modality $m$ at institution $n$
$\Gamma_{\mathcal{M}_{n},n}^{(t)}$ DOGR coefficent at global aggregation step $t$ for the classifier of institution $n$
$\varphi_{n}^{(t)}$ DOGR normalization coefficient for institution $n$ at global aggregation step $t$
~$\widetilde{\nabla}_{m}\mathcal{L}_{n}$ Stochastic gradient of institution $n$’s ML model with respect to the modality $m$ encoder parameters
~$\widetilde{\nabla}_{\bm{\varpi}}\mathcal{L}_{n}$ Stochastic gradient of institution $n$’s ML model with respect to the classifier parameters
$\Delta\bm{\omega}_{C,n}^{(t)}$ Cumulative gradient of institution $n$’s local ML model with modality combination $C$ in global aggregation step $t$
$\Delta\bm{\omega}_{C,G}^{(t)}$ Gradient of the global model with modality combination $C$ from global aggregation step $t-1$ to $t$
@@
This global model is broadcast to institution, using which they synchronize their local models as mentioned above (i.e., $\bm{\omega}^{(t+1),0}_{n}\leavevmode\nobreak\ \leftarrow\leavevmode\nobreak\ \bm{\omega}^{(t+1)},\leavevmode\nobreak\ \leavevmode\nobreak\ \forall n$) and initiate the next round of local ML training and global aggregation.
@@ III-C1 Multi-modal FL under Unbalanced Modalities
In real-world scenarios, institutions may have access to various subsets of modalities. This phenomenon makes the conventional multi-modal FL, which is based on naïve extension of uni-modal FL to multi-modal settings [31, 32, 20] impractical, as these works presume the existence of the same set of data modalities across all the institutions. In particular, in a network of institutions with non-uniform number of data modalities, the number of encoders used across them will differ. Further, the size of the input feature fed to the classifier will also vary across institutions. These variabilities make the naïve ML training and aggregation methods of [31, 32, 20] ineffective. Consequently, we propose a generalized distributed encoder-classifier architecture, which increases the flexibility of model aggregations at the server by aggregating (i) each modality’s encoder and (ii) classifiers of different existing modality combinations across the institutions separately. A schematic of our proposed system and model aggregation strategy is depicted in Fig. 1(a) and 1(b), respectively.
@@
Let $\mathcal{M}$ and $\mathcal{Y}$ denote the sets encompassing all data modalities within our network distributed across the institutions, and the data labels (i.e., cancer stages in our scenario), respectively. Moreover, let $\mathcal{M}_{n}\subseteq\mathcal{M}$ denote the subset of modalities held by institution $n$, which are present in its local dataset $\mathcal{D}_{n}$. Considering each institution $n$, each data point $d\triangleq(\bm{X},\bm{y})\in\mathcal{D}_{n}$ is associated with a feature vector $\bm{X}$ and a label $\bm{y}$, where $\bm{X}=\{\mathbf{x}_{m}\}_{m\in\mathcal{M}_{n}}$ contains the set of features corresponding to each modality $m$ possessed by the institution. Additionally, $\bm{y}\in\mathcal{Y}$ is represented in a one-hot encoded format, capturing the cancer stage. As will be formalized next, to handle institutions with an uneven distribution of modalities, we will put a greater emphasis on improving the local training of encoders. This approach diverges from the recent methodologies in multi-modal FL [32], where the emphasis lies on aggregating the classification layers – and co-attention layers – which adapt well to scenarios where institutions have access to all the modalities. This paradigm shift allows our approach to extend the distributed ML training framework of FL for each modality to all institutions possessing data pertaining to that specific modality. In particular, in our approach, the encoder used for each modality would benefit from the commonality of data across the institutions possessing that modality. This, in turn, eliminates the restricting assumption that institutions must possess all data modalities.
@@
where $\bm{\omega}_{m_{i},n}(\mathbf{x}_{m_{i}})$, for each $i\in\{1,\cdots,|\mathcal{M}_{n}|\}$, represent the encoder output upon being fed with data feature of modality $m_{i}$. Also, the function $\bm{\varpi}_{n}(\cdot)$ encompasses the description of the classifier, which takes as its input the concatenated/fused feature of all the modalities available at the institution and produces a predicted label as its output.
Fig. 2: A schematic of the encoder-classifier structure at an arbitrary institution $n$ with three data modalities.
@@ III-C2 Machine Learning Task Formulation
We consider the loss function at institution $n$ for an arbitrary local model parameter realization $\bm{\omega}_{n}$ as $\mathcal{L}_{n}(\bm{\omega}_{n};\mathcal{D}_{n})=\frac{1}{|\mathcal{D}_{n}|}\sum_{d\in\mathcal{D}_{n}}\ell(\bm{\omega}_{n};d)$ where $\ell(\bm{\omega}_{n};d)$ is the ML model loss function evaluated on datapoint $d$. In this work, we use cross-entropy to measure the loss [34]. Also, we use $\bm{\omega}_{n}\triangleq\left[\{\bm{\omega}_{m,n}\}_{m\in\mathcal{M}_{n}}:\bm{\varpi}_{n}\right]$ to denote the concatenation of the encoders model parameters in conjunction with the classifier for institution $n$.
@@
$\mathcal{P}(\mathcal{M})\triangleq\{\cup_{n\in\mathcal{N}}\mathcal{M}_{n}\}.$ (2)
If all the institutions have the same modality ($\mathcal{M}_{n}=\mathcal{M},\forall n$), then $\mathcal{P}(\mathcal{M})$${}\{\mathcal{M}\}$. Subsequently, since institutions with the same combinations of modalities use the same architecture of classifiers, we denote $\widehat{\bm{\varpi}}_{C}$ as the common architecture of the classifier used in the system, for all institutions $n$ with modality combination $C\in\mathcal{P}(\mathcal{M})$, for which $|\widehat{\bm{\varpi}}_{C}|=|\bm{\varpi}_{n}|$, $\forall n$, where |.||.|| . | captures the number of classifier parameters. [^5] For example, if the system has one institution with modality $m_{1}$ (i.e., $\mathcal{M}_{1}=\{m_{1}\}$), one institution with modality $m_{2}$, one institution with modality $m_{3}$, one institution with modality $m_{1},m_{2}$, and one institution with modality $m_{1},m_{2},m_{3}$, then, $\mathcal{P}(\mathcal{M})=\{\{m_{1}\},\{m_{2}\},\{m_{3}\},\{m_{1},m_{2}\},\{m_{1},m_{2},m_{3}\}\}$. In this case, $C$ can be any of the five sets that constitute $\mathcal{P}(\mathcal{M})$.
@@
$\mathcal{L}(\bm{\omega}_{\mathsf{obj}})\triangleq\sum_{n\in\mathcal{N}}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\mathcal{L}_{n}(\bm{\omega}_{n}),$ (3)
where $\bm{\omega}_{\mathsf{obj}}$ encompasses all the encoder and classifier parameters across the institutions (i.e., $\bm{\omega}_{\mathsf{obj}}=\{\bm{\omega}_{n}\}_{n\in\mathcal{N}}$).
Therefore, our goal is to find a set of global model parameters, consisting of a set of unified model parameters for encoders used for different modalities and a set of unified classifiers used for existing modality combinations in the network. Mathematically, we are interested in obtaining $\widehat{\bm{\omega}}_{\mathsf{obj}}^{\star}$, which is given by
$\widehat{\bm{\omega}}_{\mathsf{obj}}^{\star}=\big{\{}\{\widehat{\bm{\omega}}^{\star}_{m}\}_{m\in\mathcal{M}},\{\widehat{\bm{\varpi}}_{C}^{\star}\}_{C\in\mathcal{P}(\mathcal{M})}\big{\}}\triangleq\operatorname*{arg\,min}_{\widehat{\bm{\omega}}_{\mathsf{obj}}}\mathcal{L}(\widehat{\bm{\omega}}_{\mathsf{obj}}),$ (4)
where $\widehat{\bm{\omega}}_{\mathsf{obj}}$ is defined similarly to $\bm{\omega}_{\mathsf{obj}}$ in (3) upon deployment of unified encoder and classifier parameters across the institutions possessing the same modality combinations; mathematically $\widehat{\bm{\omega}}_{\mathsf{obj}}=\{\bm{\omega}_{n}\}_{n\in\mathcal{N}}=\{\left[\{\bm{\omega}_{m,n}\}_{m\in\mathcal{M}_{n}}:\bm{\varpi}_{n}\right]\}_{n\in\mathcal{N}}$ if $\bm{\omega}_{m,n}=\widehat{\bm{\omega}}_{m}\leavevmode\nobreak\ \forall n$ and $\bm{\varpi_{n}}=\widehat{\bm{\varpi}}_{C}$ if institution $n$ possesses modality combination $C$. In particular, $\widehat{\bm{\omega}}_{\mathsf{obj}}^{\star}$ is the set of all global optimal parameters of all encoders and classifiers present in the network, where $\widehat{\bm{\omega}}_{m}$ refers to the unified encoder parameter used for modality $m$ across all the institutions, and $\widehat{\bm{\varpi}}_{C}$ refers to the unified parameters of the classifiers used in institutions possessing the modality combination $C$ in the network. In the following, we describe the iterative procedure of local and global model aggregations, aiming to learn the optimal values of these parameters, i.e., $\{\widehat{\bm{\omega}}^{\star}_{m}\}_{m\in\mathcal{M}},\{\widehat{\bm{\varpi}}_{C}^{\star}\}_{C\in\mathcal{P}(\mathcal{M})}$. [^6] It is noteworthy to mention that upon having the same set of data modalities across the institutions, $\mathcal{P}(\mathcal{M})=\{\mathcal{M}\}$, $\widehat{\bm{\omega}}_{\mathsf{obj}}$ would contain the global encoder and classifier parameters that will be deployed uniformly across all the institutions upon the conclusion of distributed ML model training.